OpenStudy (anonymous):

What is the equation of the line that has a slope of -1 and goes through point (0, 5)?

OpenStudy (anonymous):

You want it in y = mx+b form

OpenStudy (anonymous):

\[\huge y-y _{1}=m(x-x _{1})\] It's pretty easy if you just use point - slope form.

OpenStudy (anonymous):

Where x1 and y1 are your points, and m is your slope.